what two stable compounds when combined approximately 60 min later will form a strong solution of hydrochloric acid? 26 percent if possible
1
Expert's answer
2012-09-14T09:50:51-0400
SO2Cl2 + H2O= HCl + H2SO3
CH3COCl +H2O= HCl + CH3COOH

actually
all acyl chlorides can give strong solution of HCL in water